lexxii:

ABCD is a rectangle find the length of each diagonal AC= 2(x-3) BD= x+5

lexxii:

@dude

dude:

 You can set these equal to each other

lexxii:

so x+5=2(x-3) then solve for x? then would you plug x into one of the equations?

dude:

Yes, do \(x+5=2(x-3)\)

lexxii:

okay then youd get x=11?

dude:

Yes, then you have to substitute back into each equation to find their measure

lexxii:

then the diagonal would be 16 units?

dude:

Yes

lexxii:

okay thank you

dude:

You're welcome :)